The grades in a statistics course for a particular semester were as follows:


x=A B C D F F G H

f= 136 60 34 12 29 1 3 1


Test the hypothesis at the 0.05 level of significance does thr distribution of grades follow poison distribution


1
Expert's answer
2021-09-26T18:18:28-0400

Given the grades in a statistics course for a particular semester.

To test: whether grades follow grades follow Poisson distribution at 0.05 level of significance.

The hypotheses to be tested are:

H0: The data is from Poisson distribution

H1: The data is not from Poisson distribution

We estimate parameter of Poisson distribution as

"λ= \frac{\sum fx}{\sum f} \\

λ=2.0809"

The Chi square test is given by

"χ^2 = \sum \frac{(f-e)^2}{e} \sim χ^2_{n-2, α}"

where the test statistic is calculated as follows:



The critical value of Chi square test is

"\u03c7^2_{n-2, \u03b1} = \u03c7^2_{6, 0.05} = 12.592"

Since Chi square calculated value greater than Chi square critical value, we reject the null hypothesis and can conclude that data is not from Poisson distribution.